OpenStudy (anonymous):

A rectangle has an area of 72 square units. The width of the rectangle is 9 units. The length of the rectangle is 2x + 4. What is the rectangle's length?

Parth (parthkohli):

The area of a rectangle is length times breadth. \( \color{Black}{\Rightarrow 9(2x + 4) = 72}\) \( \color{Black}{\Rightarrow 2x + 4 = 72 \div 9}\) \( \color{Black}{\Rightarrow 2x + 4 = 8 }\)

Parth (parthkohli):

I don't know why we have \(x\) here. But solving for \(x\), you get \(2x = 4\) \( \color{Black}{\Rightarrow x = 2}\)

OpenStudy (anonymous):

these are my answers 6 7 8 9

OpenStudy (anonymous):

6 7 8 9

Parth (parthkohli):

Well as I said, 2x + 4 = length = 8

OpenStudy (anonymous):

oh

OpenStudy (ganpat):

area for rect = l * b l = 2x + 4 b = 9 72 = (2x +4) * 9 x = 2 Therfore, Length = 2(2) + 4 = 8